Restart Night #11 UT 2018-09-12

OSA: J. Williams (LBTO) S. Allanson (Tucson)

Observers/ISA: B. Rothberg, J. Hill, (Tucson)

Software: S. Hooper/P. Kubanek (on-call)

ISp: E. Solheid

Mtn Manager: M. Wagner

Details

UT 01:58 - D. Miller working on vibration testing - delayed opening (RETROREF on DX)

UT 01:59 - LBCs up (no issue blue tracker) - taking test biases (high bay lights are on), LUCI software GUIs up, monitoring temps

UT 02:07 - waking MODS-2 Blue

UT 02:18 - taking MODS-2 Blue sieve snap

UT 02:23 - VFLAT did not turn on for MODS-2

UT 02:24 - taking MODS-2 imaging cals while other work is done - decided not to, D. Miller gave me the evil eye over audio

4D Data Collection on DX with quiet telescope

UT 02:26 - LUCI coldheads off, TH fans off, MV off, M1 lowered

UT 03:59 - D.Miller finished with 4D testing - powered up MCE/ROE

UT 04:40 - RetroRef will be removed now

UT 04:41 - running checklist for opening, got a LUCI-2 MOS error - failed to initizlize, pre-conditions not met. Looks like the strain gauges on the grabber are both on. Mask was in FPU before start of night.

UT 05:03 - Dave T. on phone, not strain gauges only

Looking at MOS GUI - Mask Handling Unit - initialization - stepped through with Dave all the issues, turned out it was strain gauges, zeroed them out in switchbox

Open

UT 05:16 - opening enclosure

Petr has installed a new GCS patch which prints the coords of all the wfs lenslets in tcslog and in the FITS header. (at present these are the coords before rotation, but AGW4 has rotation angle 0)

UT 05:31 ACQUIRE preset to adjust pinting.

UT 05:45 ACTIVE preset on-axis to collimate - GCS is able to converge on the initial collimation by itself. Seeing is 0.7 on wfs.

UT 05:50 ACTIVE preset off-axis x=0, y=103 runs away in coma on both sides

UT 05:55 - collimating on Stone O Field - attempting to get IDL collimation working - testing by going back and forth between on and off-axis stars

UT 06:04 - sending preset to MOS field for Stone O

UT 06:07 - preset missed on right, going to pointing star

UT 06:11 - putting masks into FPU, working on pointing which is drifting badly

UT 06:20 - Stop and restart GCSL and GCSR to revert back to version of 20180907.

UT 06:34 - continuing work on collimation with IDL

Summary of Active Optics

And now I'm speaking only of GCS 20180907, and nothing of today's patch, for the general discussion of wfs problems......

I. GCS and IDL both collimate fine for on-axis stars on the left and the right.

II. GCS and IDL both collimate poorly for typical off-axis stars (on the Y axis) on the left side only. In the limiting case, this seems not to be a GCS/IDL problem. Our first two suspects were:

a) Mid-spatial scale errors on M2 which create a spurious trefoil signal, but AdSec says the flat was loaded with 648 modes. That could imply that the flat has drifted over 6 months - something we have not observed before.

b) Perhaps this wfs camera in AGW4 has peculiar lenslets. This also seems unlikely, but something is perturbing the slopes at the edge of the pupil.

III. For GCSL only (not GCSR), the collimation runs away for some off-axis guide star positions. This appears to be cause by the pupil position being incorrectly shifted. See the example at 07:00 UT (left_wfsc iamges 00100 - 00120). The guide star is at KFP coords x=-60,y=+96 mm. Both GCS and IDL agree that the shift is -13.51 and +16.96 pixels, but GCSL reports the wrong pupil position which is erroneously always near Y=215.

Mask acquisition to measure rotator zeropoint

UT 07:00 - sending preset for Stone O

UT 07:00 - LUCI-2 MOS error - mask to turnout error

UT 07:15 - Dave T. recovered MOS error, init MOS from IMGUI

UT 07:19 - sending preset again - for some reason the mask put itself out of FPU (LUCI-2) but I don't recall sending that command

UT 07:29 - sending preset yet again - collimating and no mos errors!

UT 07:39 - LUCI-1 need to offset fieldstop because boxes are outside of image

LUCI-2 had to exclude Star #32,#36

Offset

X shift = -0.089"

Y shift = -2.200"

dtheta = 0.152

LUCI-1 is garbage will have to realign

UT 07:47 - putting LUCI-1 mask in turnout to take sky image to adjust FM4

adjust#1 2043,15 to 2048,1

adjust#2 25,1022 to 1,1022

UT 07:53 - LUCI-1 aligned, sending preset again

UT 07:54 - stopped zernikes on left

LUCI-1 exclude #34, 36, 37, 31

X-shift = 1.252"

Y-shift = 0.347"

dtheta = -0.131

error= 0.092,0.086

LUCI-2 exclude #36,37 - This one looks better aligned than earlier

X-shift = -0.216"

Y-shift = -2.144"

dtheta = 0.182

error= 0.087,0.078

LUCI Pointing Data

Load IRAF on RM507-N computer (for now)

at IRAF cl> PtStars

-enter current or about to be current LST

N30FieldStop, N375 Camera, Br Gamma, 2.51sec NDIT=2

1) acquire to first star to center up

for each star in rest of lest

2) active preset to collimate on-axis, then park probe

(first iteration of this we did active, then resent as track)

off images:

luci1.20180912.0012, luci2.20180912.0009

Sorted in Azimuth:
Catalog is PT10m
AZ ....ALT ........RA J2000 .....DEC ........... PtStarID ............L1 ..............L2
======= ===================== ===================

0.55 51.10 00:19:22.047 71:35:43.13 PT10m001..............0013.....0010
7.68 30.55 07:39:34.179 83:07:24.43 PT10m034..............0014.....0011
26.25 46.61 03:44:52.171 65:39:15.91 PT10m017..............0016.....0013
28.13 75.44 01:13:44.415 45:11:40.71 PT10m008..............0017.....0014
51.23 36.63 05:14:06.028 48:10:24.74 PT10m024..............0018.....0015
54.40 57.44 03:11:04.380 45:58:09.12 PT10m015..............0019.....0016
72.62 31.26 05:14:34.108 29:40:48.20 PT10m025..............0020.....0017
86.77 45.25 03:57:36.752 24:39:07.47 PT10m019..............0021.....0018
95.30 67.91 02:15:28.469 28:07:25.47 PT10m011..............0022.....0019
111.16 50.82 03:03:54.241 13:06:35.48 PT10m014..............0023.....0020
137.60 48.01 02:22:15.469 -0:48:41.75 PT10m012..............0025.....0021
151.22 77.73 01:00:18.152 21:47:27.61 PT10m005..............0026.....0022
152.68 31.93 02:13:14.566 -20:24:41.88 PT10m010..............0028.....0023
183.71 40.53 00:23:13.295 -16:41:32.80 PT10m003..............0029.....0024
186.39 56.26 00:20:49.768 -0:52:15.96 PT10m002..............0030.....0025
209.91 67.51 23:49:52.999 12:42:56.46 PT10m102..............0031....0026
261.84 73.35 23:19:22.350 28:54:26.91 PT10m100..............0032....0027
312.01 68.42 01:13:44.415 45:11:40.71 PT10m008..........0033........0028
127.89 46.90 04:55:39.472 2:22:11.96 PT10m022.............0034.....0029
164.76 44.73 03:28:55.041 -11:20:05.27 PT10m016..........0035.....0030
111.04 50.71 05:59:26.456 13:06:46.34 PT10m027.......0036.....0031
346.05 76.24 03:11:04.380 45:58:09.12 PT10m015.......0037.....0032

Note added later: Because of developing problems with the AGW4 guide probe, this pointing data may not be useful on the SX side.

UT 09:49 JMH changed focus coefficient left elevation lookup table to match the right side coefficients.

UT 10:03 - seeing has popped to like 1.5, clouds running around horizon inching into north, so not the best conditions

UT 11:24 - putting LUCIs in safe mode, re-configuring for LBCs (pretty pic or two? and skyflats)

UT 11:27 - Put MODS-2 to sleep

UT 11:54 - have a little time before sky brightens, ran IC342 CP that failed a few days ago, worked fine, running dohybrid

UT 12:05 - co-pointed, starting IC342 Bino 4pt script - we'll see....

UT 12:12 - stopping go straight to skyflat r-sloan, r-sloan PA-180

UT 12:14 - first r-sloans too low counts 700 -

UT 12:19 - r-sloan r-sloan (scale=3) PA=180 -first exposures 4k

UT 12:12 - taking another set PA=180 scale=1 - 10 exposures should have enough S/N for use

UT 12:30 - taking sloan-g, sloan-i PA=180 scale=0.5 (not sure this is good)

UT 12:33 - Finished Skyflats

UT 12:35 turning off LBCs - closing enclosure

-- %USERSIG{JohnHill - 2018-09-12}%
Topic revision: r14 - 12 Oct 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback